Question 12

Below are financial statements for a firm that paid a net $16 million to shareholders in 2010. Supply the missing numbers (indicated by the capital letters) in the financial statements. Make sure to show how you got to the numbers.

Balance sheet, end of 2010

Balance Sheet, End of 2010 (In Million Dollars)

2010

2009

2010

2009

Operating Assets

A

400

Operating Liabilities

200

150

Financial Assets

300

B

Financial Obligations

F

E

Shareholders’ Equity

D

180

Total

900

C

Total

G

830

Income Statement, 2010

Operating income after tax

H

Net Financial expense, after tax

4

Comprehensive income

36

B. Calculate free cash flow for 2010

C. Calculate the cash flow with respect to net financial obligations

D. Calculate return on common equity ROCE and return of net operating assets RNOA

     for 2010 on beginning of year denominators.

E. Show that ROCE and RNOA reconcile according to the financing leverage equation.
Answer #1

Missing figures

Basic formula:- Asset = Liabilities + Capital

A = 900 - 300 = 600

C = 830 (should be equal to the liabilities of 2009)

B = 830 - 400 = 430

E = 830 - 180 - 150 = 500

G = 900

Share holder's equity will be reduced by 16 million as company paid dividends to share holders in 2010.

Share holder's equity on 2009 = 180 Million

Share holder's equity on 2010 (D) = 180 - 16 = 164

F = 900 - 164 - 200 = 536